OpenStudy (jackfrost33):

please help me! The value of x in the proportion 1/2:2/3 = 3/4:x is... 1 4/9 1 7/9 1/4

OpenStudy (texaschic101):

ratios can be written as fractions. 1/2 : 2/3 can be written as (1/2) / (2/3). so set up your proportion, cross multiply, and solve for x (1/2) / (2/3) = (3/4) / x

OpenStudy (texaschic101):

do you know how to cross multiply ?

OpenStudy (jackfrost33):

yes

OpenStudy (texaschic101):

so what did you get when you cross multiplied ?

OpenStudy (jackfrost33):

495/1000 but not sure if that's correct

OpenStudy (texaschic101):

1/2x = (2/3 * 3/4) 1/2x = 6/12.....now we divide both sides by 1/2 x = (6/12) / (1/2) do you know how to divide fractions ?

OpenStudy (jackfrost33):

no i don't

OpenStudy (texaschic101):

(6/12) / (1/2) to divide fractions, you just flip what you are dividing by, then multiply 6/12 * 2/1 = 12/12 = 1 now lets check our answer.. (1/2) / (2/3) = (3/4) / (1) sub in 1 for x....and cross multiply (2/3)(3/4) = (1/2)(1) 6/12 = 1/2 1/2 = 1/2 (correct) so x = 1

OpenStudy (jackfrost33):

thx:)

OpenStudy (texaschic101):

anytime :)